Exhibit I.2 and I.3 present the association of sleep duration with any CVD stratified by age and gender, respectively. Calculate the prevalence rates for each sleep duration stratum for each age and gender category and the PRs. .

Exhibit I.2
Age < 60 Age > 60
#
Cases
#
Risk
Rate/100 PR
(vs.7 hours
#
Cases
#
Risk
Rate/100 PR
(vs.7 hours
≤ 5 hours 147 1,911 7.69 157 625 25.12
6 hours 160 5,000 3.2 258 1,369 18.85
7 hours 119 7,154 1.66 315 1,969 16
8 hours 145 6,914 2.1 494 2,774 17.81
≥ 9 hours 57 1,526 3.74 294 1,155 25.45
.
Answer #1

Prevalence rate is the total ( new and pre-existing ) number of cases of a disease that are present in a particular population at a given time.

Formula = Number of old and new cases of a disease / Total population at risk of getting that disease × 1000

For population less than 60years of age

Time less than or equal to 5 hours sleep, prevalence rate of CVD = total number of cases (new and old ) / total number of population at risk ×1000

= 147/ 1911 × 1000

= 7.69

Time is equal to 6 hours sleep, prevalence rate of CVD = total number of cases (new and old ) / total number of population at risk ×1000

= 160/ 5000×1000

= 3.2

Time is equal to 7 hours sleep, prevalence rate of CVD = total number of cases (new and old ) / total number of population at risk ×1000

= 119/ 7154

= 166

Time = 8 hours sleep, prevalence rate of CVD = total number of cases (new and old ) / total number of population at risk ×1000

= 145/6914

= 2.1

Time more than or equal to 9 hours sleep , prevalence rate of CVD = total number of cases (new and old ) / total number of population at risk ×1000

= 57/1526×1000

= 3.74

For population more than 60years of age

Time less than or equal to 5hours sleep ,prevalence rate of CVD = total number of cases (new and old ) / total number of population at risk ×1000

= 157/ 625×1000

= 25.12

Time is equal to 6 hours sleep, prevalence rate of CVD = total number of cases (new and old ) / total number of population at risk ×1000

= 258/ 1369× 1000

= 18.85

Time is equal to 7 hours sleep , prevalence rate of CVD = total number of cases (new and old ) / total number of population at risk ×1000

= 315/1969×1000

= 16

Time = 8 hours sleep, prevalence rate of CVD = total number of cases (new and old ) / total number of population at risk ×1000

= 494/ 2774× 1000

= 17.81

Time is more than or equal to 9 hours sleep , prevalence rate of CVD = total number of cases (new and old ) / total number of population at risk ×1000

= 294/ 1155 × 1000

= 25.45